OpenStudy (anonymous):

how does tan(5pi/4) = 1

OpenStudy (anonymous):

because sine and cosine are the same number there

OpenStudy (anonymous):

how do you know that?

OpenStudy (b87lar):

tan has a period of pi, so tan(x+pi) = tan(x)

OpenStudy (dannyo19):

tan(5pi/4)= sin(5pi/4)/cos(5pi/4)

OpenStudy (b87lar):

and tan(1/4 pi) = 1

OpenStudy (dannyo19):

So add a pi to 1/4pi and you're back at 1 given it's period=pi

OpenStudy (campbell_st):

well \[\tan(\frac{5\pi}{4}) = \tan(\pi + \frac{\pi}{4})\] so it's 3rd quadrant where tan is positive you can use the sum indentity to show \[\tan(\pi + \frac{\pi}{4}) = \frac{\tan(\pi) + \tan(\frac{\pi}{4})}{1 - \tan(\pi)\times \tan(\frac{\pi}{4})}\] hope that helps

OpenStudy (anonymous):

i know it because the corresponding point on the unit circle is \((\frac{-\sqrt2}{2},\frac{-\sqrt2}{2})\)

OpenStudy (anonymous):

what you need is a nice unit circle cheat sheet

OpenStudy (anonymous):

locate \(\frac{5\pi}{4}\) on the unit circle on the last page of the attached cheat sheet you will see that the first coordinate (cosine) and the second coordinate (sine) are the same there that means that the tangent is one

OpenStudy (anonymous):

I wish I could medal all of you

OpenStudy (anonymous):

Thank you

OpenStudy (campbell_st):

sorry@satellite73, didn't mean to push in